# Nonreciprocal Photon-Phonon Entanglement in Kerr-Modified Spinning Cavity Magnomechanics

### 09 Dec 2023 v1

@@ IV The nonreciprocal photon-phonon entanglement
Figure 6: The photon-phonon entanglement versus $\kappa_{m}/\kappa_{a}$ with (a) $\Delta_{K}=0$, (b) $\Delta_{K}>0$, (c) $\Delta_{K}<0$, wehre the Sagnac effect are all considered. (d) The nonreciprocity of the photon-phonon entanglement induced by the magnon Kerr effect versus $\kappa_{m}/\kappa_{a}$ with and without the Sagnac effect. Other parameters in (a-c) are the same as those in Fig. 2.
@@
In fact, the magnomechanical coupling strength $G_{mb}$ can be fine-tuned by adjusting the Rabi frequency of the driving field on the cavity [see Eq. (III.1)] in our proposal. So how does the magnomechanical coupling affect the photon-phonon entanglement with or without the magnon Kerr and Sagnac effects? To show this, we plot $E_{ab}$ versus the normalized coupling strength $G_{mb}/g_{ma}$ in Figs. 5(a-c). Obviously, the photon-phonon entanglement increases first to its maximal value and then decreases with the magnomechanical coupling strength $G_{mb}/g_{ma}$. Specifically, the Sagnac effect can only give a weak nonreciprocity on the photon-phonon entanglement in the absence of the magnon Kerr effect $\Delta_{K}=0$ [see Fig. 5(a)]. In the presence of the magnon Kerr effect ($\Delta_{K}\neq 0$), we find that a visible nonreciprocity on the photon-phonon entanglement can be induced by the Sagnac effect [see Fig. 5(b) or 5(c)]. This is because that the photon-phonon entanglement can be significantly enhanced (reduced) when $\Delta_{K}\Delta_{F}<0$ ($\Delta_{K}\Delta_{F}>0$). From Figs. 5(a-c), the magnon Kerr effect induced nonreciprocity on the photon-phonon entanglement can also be revealed with or without the Sagnac effect. More intuitively, we plot the difference ($\Delta E_{ab}$) of the logarithmic negativities between the cases of $\Delta_{K}>0$ and $\Delta_{K}<0$ in Fig. 5(d), where $\Delta E_{ab}$ is defined as
@@
When the Sagnac effect is included, we find that large $\Delta E_{ab}$ can be obtained, as shown by the red and blue curves.
Besides the magnomechanical coupling strength, the decay rate of the magnons in the Kittel mode of the YIG sphere can also be adjusted experimentally via changing the distance between the YIG sphere and the microwave attenna. We find that the optimal photon-phonon entanglement can be realized by tuning $\kappa_{m}$ when other paraters are fixed, as shown in Figs. 6(a-c). In the absence of the magnon Kerr effect, i.e., $\Delta_{K}=0$, one can see that the photon-phonon entanglement is rubust against the Sagnac effect for the small decay rate of the Kittel mode [see Fig. 6(a)], but when the magnon Kerr effect is taken into account, i.e., $\Delta_{K}\neq 0$, the visible nonreciprocity induced by the Sagnac effect can be predicted [see Fig. 6(b) or 6(c)]. We also show that the nonreciprocity induced by the Sagnac effect in the presence or absence of the magnon Kerr effect can be suppressed by increasing $\kappa_{m}$. This means that the nonreciprocity induced by the Sagnac effect can only be observed for the proper value of $\kappa_{m}$. The similar result can also be obtained for the magnon Kerr effect induced nonreciprocity of the photon-phonon entanglement with or without the Sagnac effect [see Fig. 6(d)].
@@
Figure 7: The photon-phonon entanglement versus the bath temperature $T$ with (a) $\Delta_{K}=0$, (b) $\Delta_{K}>0$, (c) $\Delta_{K}<0$, wehre the Sagnac effect are all considered. (d) The nonreciprocity of the photon-phonon entanglement induced by the magnon Kerr effect versus $T$ with and without the Sagnac effect. Other parameters in (a-c) are the same as those in Fig. 2.
@@
Next, we check the effect of the bath temperature on the photon-phonon entanglement in our proposal. For this, we plot the logarithmic negativity $E_{ab}$ versus the temperature $T$ with or without the magnon Kerr and the Sagnac effects in Fig. 7(a-c). Figure 7(a) shows that the Sagnac effect can only cause a slight enhancement on the photon-phonon entanglement at $\Delta_{K}=0$. When the magnon Kerr effect is included, we find that both the photon-phonon entanglement and its survival temperature are nonreciprocally improved (reduced) by the Sagnac effect at $\Delta_{K}\Delta_{F}<0~{}(\Delta_{K}\Delta_{F}>0)$ [see Figs. 7(b) and 7(c)]. Notably, the survival temperature of the photon-phonon entanglement in our proposal can be improved to $\sim 3$ K, which is about $15$ times than the previous proposal Li et al. (2018). Figure  7(d) also demonstrates that the magnon Kerr effect has the same effect on photon-phonon entanglement as the Sagnac effect.
@@ V CONCLUSION
In summary, we have proposed a proposal to generate a nonreciprocal photon-phonon entanglement in a Kerr-modified spinning cavity magnomechanics. The mean magnon number here can selectively display nonreciprocal linear or nonlinear (bistable) behavior with the strength of the strong driving field, where the nonreciprocity arises from the Sagnac effect, and the linear (nonlinear) behavior is the result of the interplay between the magnon Kerr effect and the magnetostrictive effect. With the enhanced magnon-phonon coupling and the swapping interaction between the magnons and the photons, magnons mediated photon-phonon entanglement is generated. This entanglement can be nonreciprocally enhanced with taking both the Sagnac and the magnon Kerr effects into account. We also show that the achieved nonreciprocal entanglement can be kept up to $\sim 3$ K with accessible parameters, exhibiting great potential in robust against the bath temperature. Our work provides a promising way to engineer various nonreciprocal devices with the magnon Kerr and the Sagnac effects in cavity magnomechanics.
ACKNOWLEDGMENTS
This work is supported by the National Natural Science Foundation of China (Grant No. 12175001 ), the Natural Science Foundation of Zhejiang Province (Grant No. Y24A040035), and the key program of the Natural Science Foundation of Anhui (Grant No. KJ2021A1301).

### 12 Apr 2024 v2

@@ IV.3 The magnomechanical coupling effect
Figure 5: The photon-phonon entanglement vs $G_{mb}/g_{ma}$ with (a) $\Delta_{K}=0$, (b) $\Delta_{K}>0$, and (c) $\Delta_{K}<0$, where the Sagnac effect is considered. (d) The nonreciprocity of the photon-phonon entanglement induced by the magnon Kerr effect vs $G_{mb}/g_{ma}$ with and without the Sagnac effect. Other parameters in panels (a)-(c) are the same as those in Fig. 3.
@@
In fact, the magnomechanical coupling strength $G_{mb}$ can be fine tuned by adjusting the amplitude of the driving field on the cavity [see Eqs. (III.1)] in our proposal. So how does the magnomechanical coupling affect the photon-phonon entanglement with or without the magnon Kerr and Sagnac effects? To show this, we plot $E_{ab}$ versus the normalized coupling strength $G_{mb}/g_{ma}$ in Figs. 5(a-c). Obviously, the photon-phonon entanglement increases first to its maximal value and then decreases to zero with the magnomechanical coupling strength $G_{mb}/g_{ma}$. Specifically, the Sagnac effect can only give a weak nonreciprocity on the photon-phonon entanglement in the absence of the magnon Kerr effect $\Delta_{K}=0$ [see Fig. 5(a)]. In the presence of the magnon Kerr effect ($\Delta_{K}\neq 0$), we find that a visible nonreciprocity on the photon-phonon entanglement can be induced by the Sagnac effect [see Fig. 5(b) or 5(c)]. This is because the photon-phonon entanglement can be significantly enhanced (reduced) when $\Delta_{K}\Delta_{F}<0$ ($\Delta_{K}\Delta_{F}>0$). From Figs. 5(a-c), the magnon Kerr effect induced nonreciprocity on the photon-phonon entanglement can also be revealed with or without the Sagnac effect. More intuitively, we plot the difference ($\Delta E_{ab}$) of the logarithmic negativities between the cases of $\Delta_{K}>0$ and $\Delta_{K}<0$ in Fig. 5(d), where $\Delta E_{ab}$ is defined as
@@
When the Sagnac effect is included, we find that large $\Delta E_{ab}$ can be obtained, as shown by the red square and blue diamond curves.
@@
Figure 6: The photon-phonon entanglement vs $\kappa_{m}/\kappa_{a}$ with (a) $\Delta_{K}=0$, (b) $\Delta_{K}>0$, and (c) $\Delta_{K}<0$, where the Sagnac effect is considered. (d) The nonreciprocity of the photon-phonon entanglement induced by the magnon Kerr effect vs $\kappa_{m}/\kappa_{a}$ with and without the Sagnac effect. Other parameters in panels (a)-(c) are the same as those in Fig. 3.
IV.4 The magnon decay rate effect
Besides the magnomechanical coupling strength, the decay rate of the magnons in the Kittel mode of the YIG sphere can also be adjusted experimentally via changing the distance between the YIG sphere and the microwave antenna. We find that the optimal photon-phonon entanglement can be realized by tuning $\kappa_{m}$ when other parameters are fixed, as shown in Figs. 6(a-c). In the absence of the magnon Kerr effect, i.e., $\Delta_{K}=0$, one can see that the photon-phonon entanglement is robust against the Sagnac effect for the small decay rate of the Kittel mode [see Fig. 6(a)], but when the magnon Kerr effect is taken into account, i.e., $\Delta_{K}\neq 0$, the visible nonreciprocity induced by the Sagnac effect can be predicted [see Fig. 6(b) or 6(c)]. We also show that the nonreciprocity induced by the Sagnac effect in the presence or absence of the magnon Kerr effect can be suppressed by increasing $\kappa_{m}$. This means that the nonreciprocity induced by the Sagnac effect can only be observed for the proper value of $\kappa_{m}$. A similar result can also be obtained for the magnon Kerr effect induced nonreciprocity of the photon-phonon entanglement with or without the Sagnac effect [see Fig. 6(d)].
IV.5 The temperature effect
Figure 7: The photon-phonon entanglement vs the bath temperature $T$ with (a) $\Delta_{K}=0$, (b) $\Delta_{K}>0$, and (c) $\Delta_{K}<0$, where the Sagnac effect is considered. (d) The nonreciprocity of the photon-phonon entanglement induced by the magnon Kerr effect versus $T$ with and without the Sagnac effect. Other parameters in panels (a)-(c) are the same as those in Fig. 3.
@@ IV.5 The temperature effect
Next, we check the effect of the bath temperature on the photon-phonon entanglement in our proposal. For this, we plot the logarithmic negativity $E_{ab}$ versus the temperature $T$ with or without the magnon Kerr or the Sagnac effect in Figs. 7(a-c). Fig. 7(a) shows that the Sagnac effect can only cause a slight improvement on both the photon-phonon entanglement and its survival temperature at $\Delta_{K}=0$. Meanwhile, we also observe the same slight improvement on both the photon-phonon entanglement and its survival temperature, as indicated by the black lines in Figs. 7(a-c) at $\Delta_{F}=0$. But when both the magnon Kerr and Sagnac effects are included, we find that the photon-phonon entanglement and its survival temperature can be significantly improved (reduced) at $\Delta_{K}\Delta_{F}<0~{}(\Delta_{K}\Delta_{F}>0)$ [see the blue line in Figs. 7(b) and the red line in Figs. 7(c)]. This indicates that the high survival temperature for the photon-phonon entanglement can only be obtained by utilizing the negative synergistic effect (i.e., the coefficients of the Kerr and Sagnac effects have opposite signs). When the positive synergistic effect (i.e., the coefficients of the Kerr and Sagnac effects have the same signs) of the Kerr and Sagnac effects is considered, the entanglement and its survival temperature reduce. Notably, the survival temperature of the photon-phonon entanglement in our proposal can be improved to $\sim 3$ K, which is about $15$ times more than the previous proposal Li et al. (2018). Figure  7(d) also demonstrates that the magnon Kerr effect has the same effect on photon-phonon entanglement as the Sagnac effect.
@@ V CONCLUSION
In summary, we have proposed to generate a nonreciprocal photon-phonon entanglement in a Kerr-modified spinning cavity magnomechanics. The mean magnon number here can selectively display nonreciprocal linear or nonlinear (bistable) behavior with the strength of the strong driving field, where the nonreciprocity arises from the Sagnac effect, and the linear (nonlinear) behavior is the result of the interplay between the magnon Kerr effect and the magnetostrictive effect. With the enhanced magnon-phonon coupling and the swapping interaction between the magnons and the photons, magnon mediated photon-phonon entanglement is generated. This entanglement can be nonreciprocally enhanced with taking both the Sagnac and the magnon Kerr effects into account. We also show that the achieved nonreciprocal entanglement can be kept up to $\sim 3$ K with accessible parameters, exhibiting great potential for robustness against the bath temperature. Our paper provides a promising way to engineer various nonreciprocal devices with the magnon Kerr and the Sagnac effects in cavity magnomechanics.
VI ACKNOWLEDGMENTS
This work is supported by the National Natural Science Foundation of China (Grant No. 12175001 ), the Natural Science Foundation of Zhejiang Province (Grant No. LY24A040004), and the key program of the Natural Science Foundation of Anhui (Grant No. KJ2021A1301).